OpenStudy (anonymous):

What is the vertex of 3x^2-6x+2

OpenStudy (anonymous):

\[ 3 x^2 - 6 x +2= 3( x^2 -2 x) +2=3(x^2 -2 x +1 -1)+2 \] Can you finish it now?

OpenStudy (radar):

a=3, b=-6, the x coordinate of the vertex =-b/2a = 6/6 =1 vertex is located at (1,y) To get y just substitute the value of 1 for x in your original formula.

OpenStudy (radar):

From a casual inspection it appears that it is at point (1,-1)